OpenStudy (anonymous):

Simply this radical equation: square root of 108 + 3 square root of 50 + 5 square root of 12 + 4 square root of 18

OpenStudy (pfenn1):

opps.....I made a mistake \[\sqrt{108}=\sqrt{2 \times 2 \times 3 \times 3 \times3}=6 \sqrt{3}\]

OpenStudy (anonymous):

no im still confused!

OpenStudy (pfenn1):

All I did was factor 108. \[\sqrt{108}=\sqrt{2\times2\times3\times3\times3}\] \[=\sqrt{2^{2}\times3^{2}\times 3} =\sqrt{2^{2}}\times \sqrt{3^{2}}\times \sqrt{3}\] \[=(2 \times 3)\sqrt{3}=6\sqrt{3}\] Do you see how I got this first term from \[\sqrt{108}\]?

OpenStudy (pfenn1):

You can factor each term of the expression this way.

OpenStudy (pfenn1):

You can factor each term of the expression this way.\[3\sqrt{50}=3\sqrt{2 \times 5 \times 5}=3\sqrt{2}\sqrt{5^{2}}=15\sqrt{2}\]